A buffer with pH = 10.15 is to be prepared by addition solid sodium hydrogen carbonate to 1.00 L of 0.300 M sodium carbonate. What mass is required assuming the volume does not change? For carbonic acid, Ka1 = 4.45 x 10-7 and Ka2 = 5.69 x 10-11
Answer #1

The balanced chemical equation is : HCO3- (aq) \rightleftharpoons CO32- (aq) + H+ (aq) : Ka2 = 5.69 x 10-11

pKa2 = -log(Ka2)

pKa2 = -log(5.69 x 10-11)

pKa2 = 10.24

According to Henderson - Hasselbalch equation,

pH = pKa2 + log([conjugate base] / [weak acid])

pH = pKa2 + log([CO32-] / [HCO3-])

10.15 = 10.24 + log(0.300 M / [HCO3-])

log(0.300 M / [HCO3-]) = 10.15 - 10.24

log(0.300 M / [HCO3-]) = -0.09

0.300 M / [HCO3-] = 10-0.09

0.300 M / [HCO3-] = 0.804

[HCO3-] = (0.300 M) / 0.804

[HCO3-] = 0.373 M

moles HCO3- = (concentration HCO3-) * (volume of solution in Liter)

moles HCO3- = (0.373 M) * (1.00 L)

moles HCO3- = 0.373 mol

moles NaHCO3 = moles HCO3-

moles NaHCO3 = 0.373 mol

mass NaHCO3 = (moles NaHCO3) * (molar mass NaHCO3)

mass NaHCO3 = (0.373 mol) * (84.0 g/mol)

mass NaHCO3 = 31.3 g